f(x) = 5x + 21/x; f(x) = -38
Answer by checkley75(3666)

f(x)=5x+21/x
-38=5x+21/x
-38=(5x^2+21)/x cross multiply
-38x=5x^2+21
5x^2+38x+21=0
(5x+3)(x+7=0
5x+3=0
5x=-3
x=-3/5 amswer.
x+7=0
x=-7 answer.
answers are (-3/5,-7).
sorry I misread the problem.